A bag of potato chips contains 595 mL of air at 30.0 ∘C and a pressure of 770 mmHg . Assuming the bag does not break, what will be its volume at the top of a mountain where the pressure is 433 mmHg and the temperature is 8.00 ∘C?
Answer #1

V=595 m2 Ti= 30°C = 30 + 273 = 303R PIE 770 mmHg Par 433 mm Hg Tz = soc= a +273= 281k Voep we know that PIVI = Pavel To 2= Pi